OpenStudy (calculusxy):

There are fewer than 40 students enrolled in a certain class. If, at a certain time, 2/9 of the students are absent from school and 1/4 of those in school have gone to the library, what is the total class enrollment? A. 18 B. 27 C. 28 D. 32 E. 36

OpenStudy (calculusxy):

@jim_thompson5910

OpenStudy (calculusxy):

@jdoe0001

jimthompson5910 (jim_thompson5910):

"2/9 of the students are absent" so 1 - (2/9) = (9/9) - (2/9) = 7/9 of the students are present agreed?

OpenStudy (calculusxy):

yes

jimthompson5910 (jim_thompson5910):

"1/4 of those in school have gone to the library" (1/4) of (7/9) = (1/4)*(7/9) = 7/36

jimthompson5910 (jim_thompson5910):

so 7/36 of the students went to the library

jimthompson5910 (jim_thompson5910):

look through the list of choices. There is only one number that when multiplied by 7/36, the result is a whole number

OpenStudy (calculusxy):

So it's E?

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (calculusxy):

Thank you! I have another three questions. Can you help me with those as well?